OpenStudy (anonymous):

S(n) is a function that squares numbers. Find S(13) when

OpenStudy (anonymous):

just did this one square thirteen

OpenStudy (anonymous):

If S(n) squares numbers, that means that S(n) = n^2. So for example, S(4)=4^2=16

OpenStudy (anonymous):